OpenStudy (anonymous):

SOMEONE PLEASE HELP! Solve the proportion x/ x^2 -2= 2/ 3x +4.. What is the value of x? x = −2 x = −2 and x = 2 x = 1 and x = 4 x = 4

OpenStudy (campbell_st):

so the problem looks like \[\frac{x}{x^2 - 2} = \frac{2}{3x + 4}\]

OpenStudy (campbell_st):

so you need to cross multiply \[x(3x + 4) = 2(x^2 - 2)\] distribute then solve for x

OpenStudy (anonymous):

i dont know how with a prob like this

OpenStudy (anonymous):

could you explain @campbell_st ?